OpenStudy (anonymous):

Gold has a density of 19.3 g/cm3. What is the mass of a 5 cm3 block of gold? A. 3.9 grams B. 19.3 grams C. 38.6 grams D. 96.5 grams

OpenStudy (anonymous):

I HAVE NO CLUE............

OpenStudy (anonymous):

mass = volume * density, You can even get that from the units. If density is 19.3 grams per cube centimeter that means that a block of 1 cm^3 has the mass of 19.3 grams. So, what is the mass of a block of 5 cm^3

OpenStudy (jamesj):

Right. By definition Mass of 1 cm^3 = Density Here Mass of 1 cm^3 = 19.3 g Hence mass of 2 cm^3 = 19.3 g + 19.3 g = 2*19.3 g mass of 3 cm^3 = 3*19.3 g etc. Hence mass of 5 cm^3 = what?

OpenStudy (anonymous):

125

OpenStudy (jamesj):

No, how did you get that?

OpenStudy (anonymous):

the 3 is cubed

OpenStudy (jamesj):

1 cm^3 = 19.3 g 2 cm^3 = 2*19.3 g = (19.3 + 19.3) g 3 cm^3 = 3*19.3 g = (19.3 + 19.3 + 19.3) g 4 cm^3 = 4*19.3 g 5 cm^3 = ... what?

OpenStudy (anonymous):

and you do that by 5

OpenStudy (anonymous):

?? confusing

OpenStudy (jamesj):

Imagine you have a little block of gold in front you. It's a cube. Each side 1 cm long. Then the volume of that cube is V = (1 cm)^3 = 1^3 cm^3 = 1 cm^3. That is, this a cube with volume of 1 cubic centimeter. The density of gold is 19.3 g/cm^3. Hence that little cube of gold weight 19.3 g. So far so good?

OpenStudy (anonymous):

yeah

OpenStudy (jamesj):

Now, suppose you have two identical such cubes in front you. What's the combined volume of those cubes?

OpenStudy (jamesj):

Two identical cubes, each with a volume of 1 cm^3. What's the combined volume of the two cubes?

OpenStudy (jamesj):

still there? talk to me or I'm moving on.